N molecules each of mass m of a gas A and 2N molecules each of mass 2m of gas B are contained in the same vessel which is maintained at temperature T. The mean square velocity of molecules of B type is v2 and the mean square rectangular component of the velocity of A type is denoted by w2. Then w2/v2


        (a)    2


        (b)    1


        (c)    1/3


        (d)    2/3


Grade:10

1 Answers

Amit Saxena
10 years ago

(d)

Mean kinetic energy of the two types of molecules should be equal. The mean square velocity of A type molecules = w2 + w2 + w2 = 3w2.

        Therefore,

        This gives m2/v2 = 2/3